# Asymptotics of the d’Arcais Numbers at Small $k$

### v1

@@ 1 Introduction
for $q=e^{2\pi i\tau}$, on the domain $\tau\in\{x+iy\,:\,x\in\mathbb{R}\,,\ y>0\}$, the upper-half-plane. This puts us in the domain of analytic number theorm. For instance, $\eta$ satisfies the modular symmetries:
@@ 3 Heuristics from the circle method
$F(t)\,=\,-\ln\big((e^{-t};e^{-t})_{\infty}\big)\,=\,\sum_{n=1}^{\infty}\frac{\sigma(n)}{n}\,t^{n}\,.$
@@
$k\ln\big(F(t_{n,k})\big)+nt_{n,k}\,=\,k\ln(\zeta(2))-k\ln(k/n)+k\,,\ \text{ as $n\to\infty$.}$
That is for the constant prefactor to the contour integral. But turning to the contour integral, we should focus on “major arcs,” and “minor arcs” of the integral.
We consider this by examining the ratio $F(t-i\theta)/F(t)$. In Figure 1, we plot $\ln(|F(t-i\theta)|^{2}/(F(t))^{2})$ and the color represents $\cos(\phi)$ for $\phi$ such that $F(t-i\theta)/F(t)=\rho e^{i\phi}$. (The the height of the curve is $\ln(\rho^{2})$.) Thus yellow represents points where the argument is nearly $\phi=0$, and blue represents points where the argument is nearly $\phi=\pi$. We have also placed black dots at all the points where $\cos(\phi)>0.99$, since these should represent the local maxima of $|F(t-i\theta)|/F(t)$ (as a function of $\theta$). On the right plot in the figure, we have plotted points $(p/q,-4\ln(q))$ for all the Farey fractions in the sequence up to denominators equal to $250$.
@@ 4 Proof of the main result
5 Some considerations for future research
A careful analysis of the remainder terms is important, and it was done for the original problem considered by Ramanujan and following researchers. Our main interest is the asymptotic behavior of $\alpha(n,k)$ as in (1). For this purpose we believe that an approach to the remainder term would be to complete the Rademacher analysis of the Hardy-Ramanujan circle method as begun in Section 3. We note that Halberstam did apply the circle method to Ramanujan’s original formula in [10].
We believe the most important open problem is to consider the $\ell=3,4,\dots$ sequence of the numbers $A(\ell,n,k)$, to see if one can obtain asymptotics for small $k$. For that, we would recommend as a starting point Abdesselam’s important recent results [2].
@@ Appendix A A detailed description of the induction step
so that $\alpha(n,k)=\rho(n,k)\beta(n,k)$. Then, by the induction hypothesis, the asymptotic equivalence statement implies the following. For each $\epsilon>0$, there is a number $N(K,\epsilon)\in\mathbb{N}$ such that we have $|\rho(n,k)-1|\leq\epsilon$ for $k=1,\dots,K$ and $n\geq N(K,\epsilon)$. Now let us use the Theorem to imply the induction step.
@@
$\alpha(m,K)\alpha(n-m,1)-(1+\widetilde{\eta}(m,K,\epsilon))(1+\widetilde{\eta}(n-m,1,\epsilon))\beta(m,K)\beta(n-m,1)\,.$

### v2

@@ 1 Introduction
for $q=e^{2\pi i\tau}$, on the domain $\tau\in\{x+iy\,:\,x\in\mathbb{R}\,,\ y>0\}$, the upper-half-plane. This puts us in the domain of analytic number theory. For instance, $\eta$ satisfies the modular symmetries:
@@ 3 Heuristics from the circle method
$F(t)\,=\,-\ln\big((e^{-t};e^{-t})_{\infty}\big)\,=\,\sum_{n=1}^{\infty}\frac{\sigma(n)}{n}\,e^{-nt}\,.$
@@
$k\ln\big(F(t_{n,k})\big)+nt_{n,k}\,\sim\,k\ln(\zeta(2))-k\ln(k/n)+k\,,\ \text{ as $n\to\infty$.}$
That is the asymptotic formula for the constant prefactor to the contour integral. But turning to the contour integral, we should focus on “major arcs,” and “minor arcs” of the integral. The major arcs are centered at every Farey fraction.
Consider the ratio $F(t-i\theta)/F(t)$. In Figure 1, we plot $\ln(|F(t-i\theta)|^{2}/(F(t))^{2})$ and the color represents $\cos(\phi)$ for $\phi$ such that $F(t-i\theta)/F(t)=\rho e^{i\phi}$. (The height of the curve is $\ln(\rho^{2})$.) Thus yellow represents points where the argument is nearly $\phi=0$, and blue represents points where the argument is nearly $\phi=\pi$. We have also placed black dots at all the points where $\cos(\phi)>0.99$, since these should represent the local maxima of $|F(t-i\theta)|/F(t)$ (as a function of $\theta$). On the right plot in the figure, we have plotted points $(p/q,-4\ln(q))$ for all the Farey fractions in the sequence up to denominators equal to $250$.
@@ 4 Proof of the main result
5 Outlook and open problems
A careful analysis of the remainder terms is important, and it was done for the original problem considered by Ramanujan and then by Ingham, by Halberstam, and by Oliver, Shrestha and Thorne. Our main interest is the asymptotic behavior of $\alpha(n,k)$ as in (1). For this purpose we believe that an approach to the remainder term would be to complete the Rademacher analysis of the Hardy-Ramanujan circle method as begun in Section 3. One advantage is that the modular symmetry is so accessible for the Dedekind-eta function.
If one wanted to apply the circle method to Theorem 2.1, then one would instead consider Eisenstein series as well as their derivatives. These also have modular or near-modular properties that are amenable to the Rademacher series analysis. We note that Halberstam did apply the circle method to Ramanujan’s original formula in [11].
We believe the most important open problem is to consider the $\ell=3,4,\dots$ sequence of the numbers $A(\ell,n,k)$, to see if one can obtain asymptotics for small $k$. For that, we would recommend as a starting point Abdesselam’s important recent results [2]. We note that his result could be considered to be a generalization of a result of Bringmann, Franke and Heim [7], in that he used a generalizations of Meinardus’ theorem. That is an approach to bypass using modularity or quasi-modularity, for example when that is not available. That is the major impediment for our elementary approach. We needed modularity to even obtain the conjecture using the first step of the Hardy-Ramanujan circle method.
@@ Appendix A A detailed description of the induction step
so that $\alpha(n,k)=\rho(n,k)\beta(n,k)$. Then, by the induction hypothesis, the asymptotic equivalence statement implies the following. For each $\epsilon>0$, there is a number $N(K,\epsilon)\in\mathbb{N}$ such that we have $|\rho(n,k)-1|\leq\epsilon$ whenever we have both $k=1,\dots,K$ and $n\geq N(K,\epsilon)$. Now let us use the Theorem to imply the induction step.
@@
$\Big(\rho(m,K)\rho(n-m,1)-\widetilde{\rho}(m,K,\epsilon)\widetilde{\rho}(n-m,1,\epsilon)\Big)\beta(m,K)\beta(n-m,1)\,.$
@@ Appendix B A review of the divisor functions
If we take a sequence of increasingly high powers of increasingly large primorials, then we can approach the limits where both inequalities are saturated. That is how we obtain the limit inferior result.